OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4 A. y – 7 = 4(x – 3) B. y – 7 = 4(x + 3) C. y + 7 = 4(x + 3) D. y + 7 = 4x – 12

OpenStudy (anonymous):

point slope form is y-y1 = m(x-x1) x1= -3 y1 = 7 and m = 4 now get the equation

OpenStudy (anonymous):

nice explaining rizwan_uet :)